Determine the electric flux through a Gaussian surface that contains 87 protons and 73 electrons.


1
Expert's answer
2022-04-18T17:31:41-0400

According to the Gauss's law, the electric flux "\Phi" through a surface is equal to:


"\Phi = \dfrac{Q}{\varepsilon_0}"

where "Q" is the total charge inside this surface, and "\varepsilon_0 = 8.85\times 10^{-12} F/m" is the vacuum permittivity. Let "e = 1.6\times 10^{-19} C" be the elementary charge. Then the total charge inside the surface is:


"Q = 87e - 73e = 14e = 22.4\times 10^{-19}C"

And the flux is:


"\Phi = \dfrac{22.4\times 10^{-19}C}{8.85\times 10^{-12} F/m} \approx 2.5\times 10^{-7}V\cdot m"

Answer. "2.5\times 10^{-7}V\cdot m".